OpenStudy (anonymous):

to find a function when do you use ≤ and < in a square root to the domain. the square root of 4-x is 4>x but for the square root of x-8 x≥8 whats the difference?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sqrt{4-x}\] domain is \(4-x\geq 0\) or \(4\geq x\)

OpenStudy (anonymous):

so when do you know what sign to use?

OpenStudy (anonymous):

@satellite73